A projectile is fired in such a way that its horizontal range is equal to three times its maximum height. What is the angle of projection?
1
Expert's answer
2021-02-19T10:32:29-0500

The expression for horizontal range of projectile is

"R=\frac{2v_0^2 sin(\theta)cos(\theta) }{g}"

The maximum height is

"H=\frac{(v_0sin(\theta))^2}{2g}"

Given that "R=3H" . So

"\frac{2v_0^2 sin(\theta)cos(\theta) }{g}=3\frac{(v_0sin(\theta))^2}{2g}"

"tan\theta =\frac{4}{2}"

So the angle of projection is "\theta =tan^{-1}(\frac{4}{3})=53^o"